OpenStudy (anonymous):

Adele is 5 years older than Timothy. In three years, Timothy will be 2/3 of Adele’s age. What is Adele’s current age? WILL MEDAL

OpenStudy (anonymous):

I believe Adele would be 6 years old, because 6 - 5 = 1, 1 + 3 = 4, and 4/6 = 2/3.

OpenStudy (anonymous):

thanks

OpenStudy (study312):

A = 5 + T T+ 3 = 2/3 ( A + 3) adele = 12 years old Timothy = 7 years old The question says in three years from now. You have to add 3 to both adele and timothy for the second equation.

OpenStudy (study312):

No direct answers on purpose here

OpenStudy (anonymous):

Oh sorry about that, I'm new :/

OpenStudy (study312):

I wasn't talking about you, I was talking about myself. No reports needed here, don't mind us

OpenStudy (anonymous):

And actually, now that I look at it I think I'm incorrect, 1 would not be 2/3 of 6

OpenStudy (study312):

2/3 of 6 is 4

OpenStudy (study312):

2 ÷ 3 = 0.6, 0.6 • 6 = 4

OpenStudy (anonymous):

Yes, but Timothy is 1, it was just that in 3 years he would be 4

OpenStudy (anonymous):

GOTTA BLAST